McCann Catching, Inc. has 2.00 million shares of stock outstanding. The stock currently sells for $12.87 per share. The firm’s debt is publicly traded and was recently quoted at 89.00% of face value. It has a total face value of $16.00 million, and it is currently priced to yield 10.00%. The risk free rate is 4.00% and the market risk premium is 8.00%. You’ve estimated that the firm has a beta of 1.49. The corporate tax rate is 38.00%.

The firm is considering a $47.38 million expansion of their production facility. The project has the same risk as the firm overall and will earn $11.00 million per year for 8.00 years.

What is the NPV of the expansion? (answer in terms of millions, so 1,000,000 would be 1.0000)
Answer #1

Equity Market Value = 2,000,000 * 12.87 = 25,740,000

Debt Market Value = 16,000,000 * 89% = 14,240,000

Toal Capital = 25,740,000 + 14,240,000 = 39,980,000

Cost of Equity = Rf + beta * Market Risk Premium

= 4% + 1.49 * 8%

= 15.92%

WACC = (Cost of Equity * Weight of Equity) + (Cost of Debt after tax * Weight of Debt)

= 15.92% * 25,740,000 / 39,980,000 + 10% * (1-0.38) * 14,240,000 / 39,980,000

= 12.46%

NPV = Present Value of Cash Inflow - Initial Outlay

= Annual Earnigns * PVAF (12.46%, 8 years) - 47,380,000

= 11,000,000 mn * 4.8880637914 - 47,380,000

= 6,396,870

NOTE: The answer to your question has been given below/above. If there is any query regarding the answer, please ask in the comment section. If you find the answer helpful, do upvote. Help us help you.